# Mechanistic Interpretability of Emotion Inference in Large Language Models

### v1

@@ Abstract
Large language models (LLMs) show promising capabilities in predicting human emotions from text. However, the mechanisms through which these models process emotional stimuli remain largely unexplored. Our study addresses this gap by investigating how autoregressive LLMs infer emotions, showing that emotion representations are functionally localized to specific regions in the model. Our evaluation includes diverse model families and sizes and is supported by robustness checks. We then show that the identified representations are psychologically plausible by drawing on cognitive appraisal theory—a well-established psychological framework positing that emotions emerge from evaluations (appraisals) of environmental stimuli. By causally intervening on construed appraisal concepts, we steer the generation and show that the outputs align with theoretical and intuitive expectations. This work highlights a novel way to causally intervene and precisely shape emotional text generation, potentially benefiting safety and alignment in sensitive affective domains.
@@
Ala N. Tak1,3,*, Amin Banayeeanzade2,3,*, Anahita Bolourani4, Mina Kian3, Robin Jia3, Jonathan Gratch1,3, 1Institute for Creative Technologies, University of Southern California (USC), 2 Information Sciences Institute, University of Southern California (USC), 3Department of Computer Science, University of Southern California (USC), 4Department of Statistics and Data Science, University of California, Los Angeles (UCLA) Correspondence: antak@ict.usc.edu, *Equal contribution
@@ 1 Introduction
Cognitive neuroscience uses functional localization approaches to identify specific brain regions responsible for particular functions and manipulate them by up/down-regulating neural activations in those regions. Akin to the shift from behaviorism to cognitive neuroscience in psychology—i.e. from treating the mind as a black box to studying brain-based cognitive processes—Mechanistic Interpretability (MI) allows for moving from black-box techniques Casper et al. (2024), to a focus on the internal mechanics of LLMs Bereska and Gavves (2024). MI can offer a fundamental understanding of how information processing is represented in LLMs, yielding fundamental insights into their inner-workings and offering new ways to control their reasoning Li et al. (2021); Rai et al. (2024); Feng et al. (2025). Building on this line of research and by drawing inspiration from emotion theory in psychology, we elucidate the inner workings of emotion processing in LLMs.
@@
In this work, we start by training linear classifiers on top of hidden representations to probe for regions where the strongest emotion-related activations occur. We provide evidence for functional localization of emotion processing and show that emotion-relevant operations are concentrated in specific layers, a consistent behavior across various model families and scales. We complement these findings by applying causal interventions, namely patching activations in the computation graphs, to identify essential components in neural representations Conmy et al. (2023); Ghandeharioun et al. (2024). As a result, we show that Multi-Head Self-Attention (MHSA) units in the mid-layers are responsible for shaping the LLM decision. To further corroborate this finding, we visualize attention patterns to show that these units consistently attend to tokens with high emotional importance. Our findings are robust and not influenced by variations in prompt wording or formatting.
@@
Additionally, we use the appraisal theory from psychology to shed light on the structure of LLMs’ internal processing. According to appraisal theory Frijda et al. (1989); Scherer et al. (1984); Smith and Ellsworth (1985), people reason about emotional situations by forming appraisal judgments (see Figure 1). We analyze the structure of emotion representations in LLMs by conducting inference-time probing on appraisal concepts to show that representations are psychologically plausible. Moreover, by modulating latent appraisal concepts to promote/demote a particular appraisal dimension (e.g., promoting self-agency), we show that the resulting changes in the output emotion align with theoretical expectations (from sadness to guilt) Marcinkevičs et al. (2024); Wu et al. (2024); Li et al. (2023).
@@
Overall, our work extends existing MI methodologies by applying them to more ecologically valid, unstructured examples, moving beyond the common practice of analyzing simplified sentence structures. Furthermore, this work serves as an early step to bridge MI techniques with applications in psychological and cognitive domains, offering insights into the inner workings of LLMs in complex, socially relevant contexts.
@@ 2 Related Work
Mechanistic Interpretability. Probing is an MI technique that uses a simple model, called a “probe”, to assess the internal representations across various layers in a model. As explained by Belinkov (2018), the groundwork for what we now refer to as probing relates back to earlier work evaluating trained classifiers on static work embeddings to predict linguistic features Köhn (2015); Gupta et al. (2015), and classified hidden states of neural models Ettinger et al. (2016); Kádár et al. (2017); Shi et al. (2016); Adi et al. (2017); Hupkes and Zuidema (2018); Belinkov (2022); Giulianelli et al. (2018). Probing is used across a variety of tasks Hewitt and Liang (2019); Tenney et al. (2019a, b); Peters et al. (2018); Clark et al. (2019); Belinkov (2018); Conneau et al. (2018).
Activation patching Heimersheim and Nanda (2024), is a causal intervention used to identify if certain activations are important to the downstream task Vig et al. (2020). By using patching, Meng et al. (2022) were able to localize where models store factual information. Patchscope, a method that extends on activation patching, is used to translate LLM representations into natural language Ghandeharioun et al. (2024).
Yet another MI technique is generation steering. This method entails manipulating a model’s activations to control the outputs Rai et al. (2024); Todd et al. (2024). Geva et al. (2022) investigated the model’s prediction process, identifying the contributions of the FFN’s output. To steer generation, they applied sub-updates promoting safety and were able to reduce the model’s toxicity. Templeton et al. (2024) find that "clamping" on features can be used to control the models’ output, steering the model’s stated goals and biases, for both desirable and undesirable outputs. Nanda et al. (2023) demonstrate that sequence models can have linear internal representations and that these representations can be used to manipulate the model’s behavior. This method closely resembles those of Turner et al. (2023) and Lieberum et al. (2023).
@@ 3 Experimental Setup
Emotion attribution is inherently subjective, making it challenging to define a single ground truth label for each input, particularly given our fine-grained list of emotions. Thus, we focus on the correctly classified examples when inspecting each language model. In other words, we only analyze the data points for which the LLM and the human annotator agreed on the same label, totaling at least 2,700 samples among different language models (see Appendix A.3 for more details). This ensures that we investigate tasks where the model performs reliably to understand the underlying mechanisms.
@@ 4 Notations and Preliminaries
Prior research suggests that both MHSA and Feed-Forward Network (FFN) units drive the generation in specific downstream tasks such as indirect object identification or concept promotion Merullo et al. (2024); Geva et al. (2022). By examining activations immediately after these units, we aim to evaluate their respective contributions to emotion processing within each transformer layer. More formally, let $\mathbf{h}^{(l)}_{t}\in\mathbb{R}^{d}$ denote the hidden state vector at layer $l$ and token index $t\in\{1,\cdots,T\}$, where $d$ is the dimensionality of the model’s hidden representations and $T$ is the input sequence length. Then,
@@
Throughout this paper, we focus on activations $\mathbf{x}_{t}^{(l)}$ selected from one of the three candidates in $\{\mathbf{a}_{t}^{(l)},\mathbf{m}_{t}^{(l)},\mathbf{h}_{t}^{(l)}\}$ and study their properties at different layers and token positions. While activations can be extracted from any layer and token, we anticipate the strongest emotion signals to be present at the last token, as it directly influences the model’s next-word prediction in a causal language modeling setup. Therefore, when clear from context, we omit the subscript $T$ while studying the last token. Also, we drop the superscript $(l)$ when generally discussing any activation across different layers.
Figure 3: Results of activation patching experiments where we measure the success rate of transferring the output emotion by patching an activation from a source sample to a target sample. The patching is performed at (Left) MHSA, (Mid) FFN, and (Right) hidden states, respectively. The MHSA and FFN heatmaps demonstrate a clear localization with successful transfer peaks happening in the mid-layers consistently across various model families and scales. This observation aligns with the consolidation points observed in the probe heatmap (Figure 2) and indicates how activation patching identifies critical regions for emotion prediction.
@@ 5 Probing for Emotion Signals
There is no clear distinction in probing performance between $\mathbf{m}^{(l)}$ and $\mathbf{h}^{(l)}$. Measurements from FFN closely track the hidden state dynamics, showing a steady increase in emotional conceptualization, peaking around the mid-layers. However, the heatmap corresponding to $\mathbf{a}^{(l)}$ shows a more dispersed pattern while following the same consistent increasing trend observed in other locations.
Our experiments reveal that emotion-processing mechanisms in LLMs are most pronounced in the middle layers across model families and sizes. Our observation aligns with the understanding that higher transformer layers capture more abstract and task-specific features. These findings suggest that the model has largely determined the output emotion by the mid-layers, with subsequent layers adding little additional processing.
@@ 6 Emotion Transfer by Activation Patching
The results shown in Figure 3 demonstrate consistent behavior across model sizes and families. When intervening on hidden state activations, we observe a clear increase in intervention success (Figure 3 right). Take Llama 3.2 1B as an example. Patching hidden states in the early layers is totally non-effective. There is a critical point, e.g. layer 10 in this model, after which copying the hidden state transfers the emotion label with a high success rate. A high success rate at the final layers is naturally an expected behavior, as the residual stream of hidden states aggregates the total information as it gets closer to the final layers. However, remarkably, the chance of success peaks and stabilizes around the mid-layers in all models, an observation that aligns with our findings in the previous section.
@@
To dig deeper, we look at the patching effect of MHSA and FFN units (see Figure 3 left and middle, respectively), which provides clear evidence for functional localization of emotion processing. More precisely, interventions targeting $\mathbf{a}^{(l)}$ and $\mathbf{m}^{(l)}$ show clear evidence of success localized to specific mid-layers, e.g., MHSA units of layers ${l\in[9-11]}$ in Llama 3.2 1B. In other words, successful patching of both MHSA and FFN units occurs only in a subset of layers and predominantly in the middle rather than the final layers, with FFN successful patching happening only slightly later than the MHSA patching. We hypothesize that there are a few consecutive layers in each language model whose MHSA units are responsible for gathering emotional information from the rest of the tokens and integrating it into the hidden state of the last token. This mechanism is immediately followed by a processing in the subsequent FFN units.
@@
To complement these findings, we perform an additional experiment, where a set of activations is knocked out in the forward pass to assess their impact. The results in Appendix B consistently align with those of our activation patching and probing, reinforcing the evidence of functional localization in emotion processing. These observations hold across different models and experimental methods, providing evidence for the universality of the identified units Rai et al. (2024).
@@
Our results in all previous sections are not prompt-dependent. Specifically, changes in the format, wording, structure, or the number of demonstrations in the prompt do not affect the findings (See Appendix C.5). Additionally, we provide a control experiment in Appendix C.4 to show that the identified units are not critical when performing a different but syntactically similar task. In fact, the final layers of the model are most critical for this syntactic task, which differs significantly from the units we found for emotion processing.
@@
To further explain our findings, we analyze the attention patterns in Llama 3.2 1B. Specifically, we record the top 3 tokens attended to by all attention heads in the last token of each layer. We conduct this analysis for all samples in the dataset, providing insight into which tokens are most frequently attended to. Figure 4 presents the aggregate attention patterns. The results indicate that early layers primarily focus on syntax. Around the mid-layers, the model shifts its attention to emotionally relevant tokens in the context and maintains this focus until the final layers. In the last few layers, attention predominantly focuses on the last token in the sequence, suggesting that it primarily carries forward the last token’s hidden state from earlier layers. These patterns provide evidence that the identified middle units are meaningfully related to emotion processing.
@@ 7 Investigating Appraisal Concepts
Figure 5 illustrates three primary appraisal dimensions and their associations with a set of basic emotions. For instance, guilt and pride are both materialized in situations with high self-agency, with the former happening in pleasant situations while the latter comes with an unpleasant experience. These mappings, which align closely with prior findings in appraisal theory Wondra and Ellsworth (2015), are extracted from our dataset by taking the average appraisal score for each emotion label.
@@
where $\hat{r}_{a}\in\mathbb{R}$ is the estimate for the appraisal score $a$, given the input $\mathbf{x}$. We train a separate appraisal probe for each layer per each appraisal at each hidden state of each layer. The weight matrices obtained in this way serve as representations of the corresponding appraisal, encoding features of the appraisal concept at the activations of the specified locations. The success of appraisal probing highly depends on whether the target concept is linearly detectable in the targeted activation. We provide the appraisal probing results in Appendix E, showing that the appraisal signals are not linearly detectable at earlier layers but are strongly present as we approach the hidden state of the final layers.

### v2

@@ Abstract
Large language models (LLMs) show promising capabilities in predicting human emotions from text. However, the mechanisms through which these models process emotional stimuli remain largely unexplored. Our study addresses this gap by investigating how autoregressive LLMs infer emotions, showing that emotion representations are functionally localized to specific regions in the model. Our evaluation includes diverse model families and sizes, and is supported by robustness checks. We then show that the identified representations are psychologically plausible by drawing on cognitive appraisal theory—a well-established psychological framework positing that emotions emerge from evaluations (appraisals) of environmental stimuli. By causally intervening on construed appraisal concepts, we steer the generation and show that the outputs align with theoretical and intuitive expectations. This work highlights a novel way to causally intervene and control emotion inference, potentially benefiting safety and alignment in sensitive affective domains. Code at: GitHub repo.
@@
Ala N. Tak1,3,*, Amin Banayeeanzade2,3,*, Anahita Bolourani4, Mina Kian3, Robin Jia3, Jonathan Gratch1,3, 1Institute for Creative Technologies, University of Southern California (USC), 2 Information Sciences Institute, University of Southern California (USC), 3Department of Computer Science, University of Southern California (USC), 4Department of Statistics and Data Science, University of California, Los Angeles (UCLA) Correspondence: antak@ict.usc.edu; banayeea@usc.edu, *Equal contribution
@@ 1 Introduction
Cognitive neuroscience uses functional localization approaches to identify specific brain regions responsible for particular functions and manipulate them by up/down-regulating neural activations in those regions. Akin to the shift from behaviorism to cognitive neuroscience in psychology—i.e. from treating the mind as a black box to studying brain-based cognitive processes—Mechanistic Interpretability (MI) allows for moving from black-box techniques Casper et al. (2024), to a focus on the internal mechanics of LLMs Bereska and Gavves (2024). MI can offer a fundamental understanding of information processing and representation in LLMs, yielding insights into their inner-workings and offering new ways to control their reasoning Li et al. (2021); Rai et al. (2024); Feng et al. (2025). Building on this line of research and by drawing inspiration from emotion theory in psychology, our work elucidates the inner workings of emotion processing in LLMs.
@@
We start by training linear classifiers on top of hidden representations to probe for regions where the strongest emotion-related activations occur. We provide evidence for functional localization of emotion processing and show that emotion-relevant operations are concentrated in specific layers, a consistent behavior across various model families and scales. We complement these findings by applying causal interventions, namely patching activations in the computation graphs, to identify essential components in neural representations Conmy et al. (2023). As a result, we show that Multi-Head Self-Attention (MHSA) units in the mid-layers are predominantly responsible for shaping the LLM decision. To further corroborate this, we visualize attention patterns, revealing that MHSA units consistently attend to emotionally loaded tokens. Our findings are robust and not influenced by variations in prompt wording or formatting.
@@
Additionally, we use the appraisal theory from psychology to shed light on the structure of LLMs’ internal representations. According to appraisal theory Frijda et al. (1989); Scherer et al. (1984); Smith and Ellsworth (1985), people reason about emotional situations by forming appraisal judgments (see Figure 1 as an example). We analyze the structure of emotion representations in LLMs by conducting inference-time probing on appraisal concepts to show that representations are psychologically plausible. Moreover, by modulating latent appraisal concepts to promote/demote a particular appraisal dimension (e.g., promoting self-agency), we show that the resulting changes in the output emotion align with theoretical expectations (from guilt to sadness) Marcinkevičs et al. (2024); Wu et al. (2024); Li et al. (2023).
@@
Overall, our work extends existing MI methodologies by applying them to more ecologically valid, unstructured examples, moving beyond the common practice of analyzing simplified sentence structures Wang et al. (2023a); Hanna et al. (2023). Furthermore, this work serves as an early step to bridge MI techniques with applications in psychological and cognitive domains, offering insights into the inner workings of LLMs in complex, socially relevant contexts.
@@ 2 Related Work
Cognitive-Neuroscientific Alignment. There is a growing body of evidence demonstrating that neural network activations can reflect cognitive constructs traditionally studied in psychology and neuroscience. For example, transformer self-attention mechanisms are shown to correlate with human eye-tracking data during reading, suggesting that LLMs may learn attention patterns that reflect human cognitive attention Bensemann et al. (2022); Eberle et al. (2022). LLMs also exhibit strong representational alignment with neural activity in the human language systems, indicating convergence between learned representations and brain-like processing Aw et al. (2024), and word embeddings show alignment with fMRI activity during word association tasks Kwon et al. (2024). Recent work even uses LLM-derived representations to predict neural activation patterns tied to high-level semantic categories like faces or places, offering scalable proxies for annotation in cognitive experiments Liu et al. (2025). Other studies localize LLM units that align with cortical language areas AlKhamissi et al. (2025) and reveal functional specialization patterns predictive of regional brain activity Kumar et al. (2024).
These findings support the plausibility of mapping psychological or cognitive theories—such as appraisal theory—onto the internal structure of LLMs. Our work builds on this foundation by examining how appraisal mechanisms are internally encoded, bridging prior behavioral studies of appraisal-emotion mappings in LLMs Tak and Gratch (2024, 2023); Broekens et al. (2023); Yongsatianchot et al. (2023) with MI.
Mechanistic Interpretability. Probing is an MI technique that uses a simple model, called a “probe”, to assess the internal representations across various layers in a model. As explained by Belinkov (2018), the groundwork for what we now refer to as probing relates back to earlier work evaluating trained classifiers on static word embeddings to predict linguistic features Köhn (2015); Gupta et al. (2015), and classified hidden states of neural models Ettinger et al. (2016); Kádár et al. (2017); Shi et al. (2016); Adi et al. (2017); Hupkes and Zuidema (2018); Belinkov (2022); Giulianelli et al. (2018). Probing is used across a variety of tasks Hewitt and Liang (2019); Tenney et al. (2019a, b); Peters et al. (2018); Clark et al. (2019); Belinkov (2018); Conneau et al. (2018).
Activation patching Heimersheim and Nanda (2024), is a causal intervention used to identify if certain activations are important to the downstream task Vig et al. (2020). By using patching, Meng et al. (2022) are able to localize where models store factual information. Patchscope, a method that extends on activation patching, is used to translate LLM representations into natural language Ghandeharioun et al. (2024).
Yet another MI technique is generation steering. This method entails manipulating a model’s activations to control the outputs Rai et al. (2024); Todd et al. (2024). Geva et al. (2022) highlight the role of Feed-Forward Network (FFN) units in promoting concepts. To steer generation, they apply sub-updates promoting safety and are able to reduce the model’s toxicity. Templeton et al. (2024) find that "clamping" on features can be used to control the models’ output, steering the model’s stated goals and biases for both desirable and undesirable outputs. Nanda et al. (2023) demonstrate that sequence models can have linear internal representations and that these representations can be used to manipulate the model’s behavior. This method closely resembles those of Turner et al. (2023) and Lieberum et al. (2023).
@@ 3 Experimental Setup
Emotion attribution is inherently subjective, making it challenging to define a single ground truth label for each input, particularly given our fine-grained list of emotions. Thus, we focus on the correctly classified examples when inspecting each language model. In other words, we only analyze the samples for which the LLM and the human annotator agreed on the same label, totaling at least 2,700 samples among different language models (see Appendix A.3 for more details). This ensures that we assess tasks where the model performs reliably to understand the underlying mechanisms.
@@ 4 Notations and Preliminaries
Prior research suggests that both MHSA and FFN units drive the generation in specific downstream tasks such as indirect object identification or concept promotion Merullo et al. (2024); Geva et al. (2022). By examining activations immediately after these units, we aim to evaluate their respective contributions to emotion processing within each transformer layer. More formally, let $\mathbf{h}^{(l)}_{t}\in\mathbb{R}^{d}$ denote the hidden state vector at layer $l$ and token index $t\in\{1,\cdots,T\}$, where $d$ is the dimensionality of the model’s hidden representations and $T$ is the input sequence length. Then,
@@
Throughout this paper, we focus on activations $\mathbf{x}_{t}^{(l)}$ selected from one of the three candidates in $\{\mathbf{a}_{t}^{(l)},\mathbf{m}_{t}^{(l)},\mathbf{h}_{t}^{(l)}\}$ and study their properties at different layers and token positions. While activations can be extracted from any token index, we anticipate the strongest emotion signals to be present at the last token, as it directly influences the model’s output prediction in our task setup. Therefore, when clear from context, we omit the subscript $T$ while studying the last token. Also, we drop the superscript $(l)$ when generally discussing any activation across different layers.
Figure 3: Results of activation patching experiments where we measure the success rate of transferring the output emotion by patching an activation from a source sample to a target sample. The patching is performed at (Left) MHSA, (Mid) FFN, and (Right) hidden states, respectively. The MHSA and FFN heatmaps demonstrate a clear localization with successful transfer peaks happening in the mid-layers consistently across various model families and scales. This observation aligns with the consolidation points observed in the probe heatmap (Figure 2) and indicates how activation patching can identify critical regions for emotion prediction.
@@ 5 Probing for Emotion Signals
There is no clear distinction in probing performance between $\mathbf{m}^{(l)}$ and $\mathbf{h}^{(l)}$. Measurements from FFN closely track the hidden state dynamics, showing a steady increase in emotional conceptualization that peaks around the mid-layers. However, the heatmap corresponding to $\mathbf{a}^{(l)}$ shows a more dispersed pattern while following the same consistent increasing trend observed in other locations.
Our experiments reveal that emotion-processing mechanisms in LLMs are most pronounced in the middle layers across model families and sizes. These findings suggest that the model has largely determined the output emotion by the mid-layers, with subsequent layers adding little additional processing. Our observation aligns with the understanding that higher transformer layers capture more abstract and task-specific features.
@@ 6 Emotion Transfer by Activation Patching
The results shown in Figure 3 demonstrate consistent behavior across model sizes and families. When intervening on hidden state activations, we observe a clear increase in intervention success (see Figure 3 right). Take Llama 3.2 1B as an example. Patching hidden states in the early layers is entirely ineffective. There is a critical point, e.g. layer 10 in this model, after which copying the hidden state transfers the emotion label with a high success rate. A high success rate at the final layers is naturally an expected behavior, as the residual stream of hidden states aggregates the information as it gets closer to the final layers. However, remarkably, the chance of success peaks and stabilizes around the mid-layers in all models, an observation that aligns with our findings in the previous section.
@@
To dig deeper, we look at the patching effect of MHSA and FFN units (see Figure 3 left and middle, respectively), which provides clear evidence for functional localization of emotion processing. More precisely, interventions targeting $\mathbf{a}^{(l)}$ and $\mathbf{m}^{(l)}$ show clear evidence of success localized to specific mid-layers, e.g., MHSA units of layers ${l\in[9-11]}$ in Llama 3.2 1B. In other words, successful patching of both MHSA and FFN units occurs only in a subset of layers and predominantly in the middle rather than the final layers, with the FFN’s most successful patchings happening only slightly later than the MHSA’s patching. We hypothesize that there are a few consecutive layers in each language model whose MHSA units are responsible for gathering emotional information from the rest of the tokens and integrating it into the hidden state of the last token. This mechanism is immediately followed by a processing in the subsequent FFN units.
@@
To complement these findings, we perform an additional experiment, where a set of activations is knocked out in the forward pass to assess their impact. The results in Appendix B consistently align with those of our activation patching and probing, reinforcing the evidence of functional localization in emotion processing. These observations hold across different models and experimental methods.
@@
Our results in all previous sections are not prompt-dependent. Specifically, changes in the format, wording, structure, or the number of demonstrations in the prompt do not affect the findings (see Appendix C.5). Additionally, we provide a control experiment in Appendix C.4 to show that the identified units are not critical when performing a different yet syntactically similar task. In fact, the final layers are most critical for this isomorphic syntactic task, which differs significantly from the units we found for emotion processing.
@@
To further explain our findings, we analyze the attention patterns in Llama 3.2 1B. Specifically, we record the top 3 tokens attended to by all attention heads in the last token of each layer. We conduct this analysis for all samples in the dataset, providing insight into which tokens are most frequently attended to. Figure 4 presents the aggregate attention patterns. The results indicate that early layers primarily focus on syntax. Around the mid-layers, the model shifts its attention to emotionally relevant tokens in the query and maintains this focus until the final layers. In the last few layers, attention predominantly focuses on the last token in the sequence, suggesting that it primarily carries forward the last token’s hidden state from earlier layers. These patterns provide evidence that the identified middle units are meaningfully related to emotion processing.
@@ 7 Investigating Appraisal Concepts
Figure 5 illustrates three primary appraisal dimensions, including pleasantness, other-agency, and predictability, in addition to their associations with a set of basic emotions. For instance, guilt and pride are both materialized in situations with high self-agency, with the former happening in pleasant situations while the latter comes with an unpleasant experience. These mappings, which align closely with prior findings in appraisal theory Wondra and Ellsworth (2015), are extracted from our dataset by averaging appraisal scores for each emotion label.
@@
where $\hat{r}_{a}\in\mathbb{R}$ is the estimate for the appraisal score $a$, given the input $\mathbf{x}$. We train a separate appraisal probe for each appraisal at the hidden states of each layer. The weight matrices obtained through this process serve as representations of the corresponding appraisal, encoding features of the appraisal concept at the activations of the specified locations. The success of appraisal probing highly depends on whether the target concept is linearly detectable in the targeted activation. We provide the appraisal probing results in Appendix E, showing that the appraisal signals are not linearly detectable at earlier layers but are strongly present as we approach the hidden state of the final layers.